# Struct2D: A Perception-Guided Framework for Spatial Reasoning in Large Multimodal Models

### v1

@@ Abstract
Unlocking spatial reasoning in Large Multimodal Models (LMMs) is crucial for enabling intelligent interaction with 3D environments. While prior efforts often rely on explicit 3D inputs or specialized model architectures, we ask: can LMMs reason about 3D space using only structured 2D representations derived from perception? We introduce Struct2D, a perception-guided prompting framework that combines bird’s-eye-view (BEV) images with object marks and object-centric metadata, optionally incorporating egocentric keyframes when needed. Using Struct2D, we conduct an in-depth zero-shot analysis of closed-source LMMs (e.g., GPT-o3) and find that they exhibit surprisingly strong spatial reasoning abilities when provided with structured 2D inputs, effectively handling tasks such as relative direction estimation and route planning. Building on these insights, we construct Struct2D-Set, a large-scale instruction tuning dataset with 200K fine-grained QA pairs across eight spatial reasoning categories, generated automatically from 3D indoor scenes. We fine-tune an open-source LMM (Qwen2.5VL) on Struct2D-Set, achieving competitive performance on multiple benchmarks, including 3D question answering, dense captioning, and object grounding. Our approach demonstrates that structured 2D inputs can effectively bridge perception and language reasoning in LMMs-without requiring explicit 3D representations as input. We will release both our code and dataset to support future research.
Figure 1: Overview of our Struct2D framework for enabling spatial reasoning in large multimodal models (LMMs). From an RGB-D video, we generate structured 2D inputs—BEV images with filtered object marks, object-centric metadata, and optional keyframes—via a 3D perception module. These inputs prompt an LMM with spatial priors and visual context, enabling diverse spatial reasoning tasks without explicit 3D input at inference.
@@ 1 Introduction
Understanding objects and their spatial relationships in 3D space is a cornerstone of intelligent interaction in complex physical environments. Tasks such as robotic manipulation [66, 37], autonomous navigation [55, 25], and visual reasoning [3, 81, 52, 14, 10, 100] all depend on accurate spatial understanding of scenes. At the core of these tasks lies the ability to localize objects precisely and reason about their configurations in 3D space. Moreover, grounding such spatial understanding in natural language enhances an AI system’s ability to interpret, explain, and act upon spatial information in human-centric contexts.
@@
Traditional task-specific models rely on explicit 3D representations as input, such as point clouds or reconstructed environments [3, 52, 33, 101], providing detailed geometric information. However, these models, often trained on limited data sources, making them less adaptable and struggle to generalize to diverse and complex textual queries. As a result, they fail to effectively bridge spatial reasoning with language comprehension, limiting their applicability for embodied AI.
@@
In recent years, Large Multimodal Models (LMMs) [95, 59, 24, 39] developed with Large Language Models (LLMs) have achieved significant advances in perception and reasoning tasks for images and videos. To extend LMMs’ capabilities to 3D understanding, point cloud-based LLMs [26, 61, 60, 78, 13, 23, 27, 53, 72, 30] have emerged, incorporating 3D spatial features by aligning point cloud data with LLMs. This integration enhances spatial reasoning and provides a richer understanding of the 3D physical world. However, they often rely on well-annotated datasets for instruction tuning and require point-cloud features as input, which limits their flexibility.
@@
Unlike models that take explicit 3D representations as input, humans perceive the world as a continuous stream of 2D visual inputs akin to a video, and naturally infer spatial relationships and object configurations by building mental representations subconsciously [57, 69]. Naturally, we ask “Can LMMs perform spatial reasoning without using explicit 3D features as direct inputs?” Recent work has begun to explore this direction by leveraging cognitive maps [81] and Bird’s Eye View (BEV) images [62] generated from video as 2D spatial cues, enabling LMMs to perform spatial reasoning [80, 84]. While promising, these approaches often omit object appearance and detailed priors (e.g., coordinates, categories), which are critical for comprehensive 3D understanding.
@@
We conduct an in-depth analysis of LMMs’ spatial reasoning abilities using a perception-guided 2D framework called Struct2D Prompting. This strategy transforms 3D perception outputs—obtained from off-the-shelf detectors—into structured 2D inputs, consisting of (1) a rendered bird’s-eye-view (BEV) image with projected object marks [^1] We follow the term “mark” as used in [80]. and (2) object-centric metadata such as category labels and 3D coordinates. When appearance cues are needed, we optionally incorporate egocentric keyframes selected based on object visibility. This design enables LMMs to reason about complex 3D scenes using only structured 2D visual and textual cues, eliminating the need for explicit 3D inputs. We begin by evaluating this approach on GPT-o3 [59], a representative closed-source LMM, to assess its zero-shot spatial reasoning capabilities.
@@
To better understand the spatial reasoning capabilities of existing LMMs, we begin with a zero-shot analysis using our proposed Struct2D Prompting strategy. The goal is to evaluate whether a pretrained, closed-source model such as GPT-o3 can accurately infer 3D spatial relationships when given only structured 2D visual and textual inputs. We use rendered bird’s-eye-view (BEV) images with projected object marks and object-centric metadata, allowing the model to reason about 3D scenes without access to explicit 3D features. This analysis yields several key insights: (1) A single informative BEV image, combined with metadata, is often sufficient for accurate zero-shot 3D scene understanding; (2) Prompt composition is critical—different spatial reasoning tasks benefit from tailored input formats; (3) For challenging tasks in VSI-Bench [81], such as egocentric-to-allocentric transformations, LMMs can perform robustly when provided with well-structured 2D projections of the 3D scene.
@@
Guided by the findings from our zero-shot analysis, we construct a large-scale instructional tuning dataset, named Struct2D-Set, using an automated pipeline. The dataset consists of 200K QA pairs generated from 6K 3D indoor scenes, leveraging ground-truth object annotations provided by the original 3D datasets. It spans eight categories of spatial reasoning tasks relevant to embodied AI. To ensure data quality, we use ChatGPT to both enrich the QA pairs with step-by-step reasoning traces and identify potentially low-quality samples. Additionally, we incorporate a human-in-the-loop review process to further refine and validate the dataset. We then fine-tune an open-source LMM (Qwen2.5VL [71]) using Struct2D-Set. Although the fine-tuned model is evaluated under noisy 3D perception conditions, it achieves strong performance across multiple spatial reasoning benchmarks, including 3D question answering [81, 3, 52], spatial captioning [14], and object grounding [10, 94], demonstrating the practicality and robustness of our approach.
@@
  • We propose a perception-guided 2D prompting strategy, Struct2D Prompting, and conduct a detailed zero-shot analysis that reveals LMMs’ ability to perform 3D spatial reasoning from structured 2D inputs alone.
@@
  • We fine-tune an open-source LMM to achieve competitive performance across several spatial reasoning benchmarks, validating the real-world applicability of our framework.
@@ 2 Related Work
3D Spatial Reasoning with MLLMs. Developing real-world embodied agents requires equipping Large Multimodal Models (LMMs) with robust 3D spatial reasoning abilities [8, 9, 16, 41, 47, 82, 98]. Recent efforts have explored spatial understanding through language [56, 64, 76], static 2D images [63, 68, 80, 54, 51], or videos [62, 81, 44]. Our work builds upon the video-input setting, but diverges by enabling spatial reasoning in LMMs using only structured 2D inputs—BEV images, object marks, and metadata—without relying on explicit 3D encoder / representations at inference.
@@
Instruction Tuning for 3D Spatial Reasoning. Recent work [34, 42, 43, 13, 44] has explored instruction tuning to enhance LMMs’ capabilities for 3D spatial reasoning, targeting tasks such as 3D visual grounding [2, 10, 94], 3D dense captioning [14], and 3D question answering [3, 52]. M3DBench [42] provides region- and scene-level instruction-response pairs for general 3D understanding, while 3DMIT [43] focuses on scene-centric instructions. LL3DA [13] supports interactive planning and reasoning across omni-3D inputs. Robin3D [34] introduces a 3D LLM trained on diverse instruction-following examples. R1-Zero-VSI [44] proposes a video-based instruction tuning dataset and a GRPO-based training method to enhance spatial reasoning in Qwen2-VL, yet its QA pairs involve limited reasoning complexity, cover fewer task types, and yield marginal performance gains. In contrast, we propose Struct2D-Set, a large-scale dataset that enables open-source LMMs to acquire rich 3D spatial reasoning skills through instruction tuning—using only structured 2D representations, without requiring direct access to 3D point clouds.
@@
3D Point Cloud LLMs. Recent advances in 3D point cloud LLMs enable natural language generation and interaction grounded in 3D geometry by directly processing point clouds as input. These models benefit from the geometric precision and texture richness of point clouds, offering stronger spatial understanding than raw image or video inputs. Prior work has focused on object-level [26, 61, 60, 78] and scene-level [13, 23, 27, 53, 72, 30] spatial reasoning. However, directly using point cloud features requires additional training and infrastructure, limiting flexibility and scalability in real-world applications.
@@
Prompting LLMs. Despite the rapid scaling of large language models (LLMs)[5, 17, 70, 92, 1, 32, 22], their reasoning capabilities remain heavily dependent on effective prompt design. In-context learning[5, 21], which conditions models on a few representative examples, has become a widely adopted technique for improving instruction-following behavior. To further enhance reasoning, strategies such as chain-of-thought [73] and tree-of-thought [85] prompting have been proposed. More recently, large multimodal models (LMMs)[59, 24, 49, 99, 91, 20, 40, 15, 79, 45, 90, 95, 39] have gained prominence for their ability to reason over multiple input modalities. This has led to a surge of research into prompting techniques tailored for LMMs[84, 80, 16, 7, 29, 58, 75, 77, 88, 28, 46, 38, 93, 97, 48, 67, 74, 83]. Building on this direction, we propose Struct2D, a structured 2D prompting strategy that enables LMMs to perform 3D spatial reasoning effectively—without requiring explicit 3D input representations.
@@ 3.1 Struct2D Prompting
Given a video $\mathbf{V}$ as input, an LMM $\mathcal{F}$ processes a set of $N$ sampled video frames, denoted as $\mathbf{I}=\{I_{1},I_{2},\dots,I_{N}\}$, where each frame $I_{n}$ has dimensions $\mathbb{R}^{H\times W\times 3}$ for $n\in\{1,...,N\}$. Alongside visual input, the LMM receives a text query of length $l_{i}$, represented as $\mathbf{T}^{\text{in}}=[t^{i}_{1},\dots,t^{i}_{l_{i}}]$. The model then generates a textual response of length $l_{o}$, denoted as $\mathbf{T}^{\text{out}}=[t^{o}_{1},\dots,t^{o}_{l_{o}}]$, formulated as:
@@
However, directly using video frames for spatial reasoning introduces two major limitations: (1) Incomplete perception — Video frames are typically sampled sparsely and from limited viewpoints, which can result in missing critical visual evidence required for spatial understanding. For instance, consider a scene where a chair is tucked partially under a table. If most sampled frames are taken from frontal views or from a standing height, the chair’s presence might be obscured or entirely invisible, leading the model to incorrectly assume there is empty space beneath the table. This limitation becomes more severe in cluttered or occluded environments, where small objects or those blocked by other furniture. (2) Lack of global context — Video frames offer fragmented, egocentric views that often fail to capture the overall structure of the scene. For example, determining whether a lamp is closer to the couch or the bookshelf may be impossible if the two objects never co-occur in the same frame. Without a consistent top-down or holistic representation, the model must rely on spatial memory or reasoning across disjoint perspectives—an ability that remains weak in most LMMs. This fragmentation also impedes the understanding of traversability (e.g., identifying a clear path from the door to the kitchen) or relational queries (e.g., which chair is directly behind the dining table).
@@
This formulation illustrates how Struct2D leverages 3D perception as an intermediate step to generate informative 2D inputs that preserve spatial structure. Although 3D point clouds are used during preprocessing, they are not directly provided to the LMM. Instead, they are transformed into BEV images and metadata used for prompting. As a result, the model performs spatial reasoning effectively without requiring explicit 3D representations as input.
@@
Evaluation Setup. We sample questions from VSI-Bench [81], which is designed to evaluate complex spatial reasoning skills. Compared to traditional 3D QA datasets [3, 52], VSI-Bench covers more fine-grained object perception requirements, intricate global spatial relationships, and egocentric-to-allocentric transformations. It also features diverse indoor scene sources and robust evaluation metrics that go beyond rule-based NLP scoring.
Comparison to GPT4Scene Prompting [62]. While GPT4Scene pioneered 2D spatial prompting using BEV images, our Struct2D strategy introduces several key improvements: (1) Filtered object marks tailored to the query improve visual relevance and reduce distraction; (2) Guided metadata prompts provide additional spatial priors; (3) Keyframe selection is optimized using depth-aware 3D projection instead of uniform sampling, making them both fewer and more informative.

### v3

# Struct2D: A Perception-Guided Framework for Spatial Reasoning in MLLMs

@@ Abstract
Unlocking spatial reasoning in Multimodal Large Language Models (MLLMs) is crucial for enabling intelligent interaction with 3D environments. While prior efforts often rely on explicit 3D inputs or specialized model architectures, we ask: can MLLMs reason about 3D space using only structured 2D representations derived from perception? We introduce Struct2D, a perception-guided prompting framework that combines bird’s-eye-view (BEV) images with object marks and object-centric metadata, optionally incorporating egocentric keyframes when needed. Using Struct2D, we conduct an in-depth zero-shot analysis of closed-source MLLMs (e.g., GPT-o3) and find that they exhibit surprisingly strong spatial reasoning abilities when provided with structured 2D inputs, effectively handling tasks such as relative direction estimation and route planning. Building on these insights, we construct Struct2D-Set, a large-scale instruction tuning dataset with 200K fine-grained QA pairs across eight spatial reasoning categories, generated automatically from 3D indoor scenes. We fine-tune an open-source MLLM (Qwen2.5VL) on Struct2D-Set, achieving competitive performance on multiple benchmarks, including 3D question answering, dense captioning, and object grounding. Our approach demonstrates that structured 2D inputs can effectively bridge perception and language reasoning in MLLMs-without requiring explicit 3D representations as input. We will release both our code and dataset to support future research.
Figure 1: Overview of our Struct2D framework for enabling spatial reasoning in Multimodal Large Language Models (MLLMs). From an RGB-D video, we generate structured 2D inputs—BEV images with filtered object marks, object-centric metadata, and optional keyframes—via a 3D perception module. These inputs prompt an MLLM with spatial priors and visual context, enabling diverse spatial reasoning tasks without explicit 3D input at inference.
@@ 1 Introduction
Understanding objects and their spatial relationships in 3D space is a cornerstone of intelligent interaction in complex physical environments. Tasks such as robotic manipulation [sisbot2007spatial, landsiedel2017review], autonomous navigation [marza2022teaching, gu2022vision], and visual reasoning [azuma2022scanqa, yang2024thinking, ma2022sqa3d, chen2021scan2cap, chen2020scanrefer, zhu2024towards] all depend on accurate spatial understanding of scenes. At the core of these tasks lies the ability to localize objects precisely and reason about their configurations in 3D space. Moreover, grounding such spatial understanding in natural language enhances an AI system’s ability to interpret, explain, and act upon spatial information in human-centric contexts.
@@
Traditional task-specific models rely on explicit 3D representations as input, such as point clouds or reconstructed environments [azuma2022scanqa, ma2022sqa3d, jin2023context, zhu20233d], providing detailed geometric information. However, these models, often trained on limited data sources, making them less adaptable and struggle to generalize to diverse and complex textual queries. As a result, they fail to effectively bridge spatial reasoning with language comprehension, limiting their applicability for embodied AI.
@@
In recent years, Multimodal Large Language Models (MLLMs) [zhang2024llavanext-video, o3, gemini, li2024llava] developed with Large Language Models (LLMs) have achieved significant advances in perception and reasoning tasks for images and videos. To extend MLLMs’ capabilities to 3D understanding, point cloud-based LLMs [guo2023point, qi2024shapellm, qi2024gpt4point, xu2024pointllm, chen2024ll3da, fu2024scene, hong20233d, man2024lexicon3d, wang2023chat, huang2023chat] have emerged, incorporating 3D spatial features by aligning point cloud data with LLMs. This integration enhances spatial reasoning and provides a richer understanding of the 3D physical world. However, they often rely on well-annotated datasets for instruction tuning and require point-cloud features as input, which limits their flexibility.
@@
Unlike models that take explicit 3D representations as input, humans perceive the world as a continuous stream of 2D visual inputs akin to a video, and naturally infer spatial relationships and object configurations by building mental representations subconsciously [nadel2008hippocampus, tolman1948cognitive]. Naturally, we ask “Can MLLMs perform spatial reasoning without using explicit 3D features as direct inputs?” Recent work has begun to explore this direction by leveraging cognitive maps [yang2024thinking] and Bird’s Eye View (BEV) images [qi2025gpt4scene] generated from video as 2D spatial cues, enabling MLLMs to perform spatial reasoning [yang2023set, yang2023dawn]. While promising, these approaches often omit object appearance and detailed priors (e.g., coordinates, categories), which are critical for comprehensive 3D understanding.
@@
We conduct an in-depth analysis of MLLMs’ spatial reasoning abilities using a perception-guided 2D framework called Struct2D Prompting. This strategy transforms 3D perception outputs—obtained from off-the-shelf detectors—into structured 2D inputs, consisting of (1) a rendered bird’s-eye-view (BEV) image with projected object marks [^1] We follow the term “mark” as used in [yang2023set]. and (2) object-centric metadata such as category labels and 3D coordinates. When appearance cues are needed, we optionally incorporate egocentric keyframes selected based on object visibility. This design enables MLLMs to reason about complex 3D scenes using only structured 2D visual and textual cues, eliminating the need for explicit 3D inputs. We begin by evaluating this approach on GPT-o3 [o3], a representative closed-source MLLM, to assess its zero-shot spatial reasoning capabilities.
@@
To better understand the spatial reasoning capabilities of existing MLLMs, we begin with a zero-shot analysis using our proposed Struct2D Prompting strategy. The goal is to evaluate whether a pretrained, closed-source model such as GPT-o3 can accurately infer 3D spatial relationships when given only structured 2D visual and textual inputs. We use rendered bird’s-eye-view (BEV) images with projected object marks and object-centric metadata, allowing the model to reason about 3D scenes without access to explicit 3D features. This analysis yields several key insights: (1) A single informative BEV image, combined with metadata, is often sufficient for accurate zero-shot 3D scene understanding; (2) Prompt composition is critical—different spatial reasoning tasks benefit from tailored input formats; (3) For challenging tasks in VSI-Bench [yang2024thinking], such as egocentric-to-allocentric transformations, MLLMs can perform robustly when provided with well-structured 2D projections of the 3D scene.
@@
Guided by the findings from our zero-shot analysis, we construct a large-scale instructional tuning dataset, named Struct2D-Set, using an automated pipeline. The dataset consists of 200K QA pairs generated from 6K 3D indoor scenes, leveraging ground-truth object annotations provided by the original 3D datasets. It spans eight categories of spatial reasoning tasks relevant to embodied AI. To ensure data quality, we use ChatGPT to both enrich the QA pairs with step-by-step reasoning traces and identify potentially low-quality samples. Additionally, we incorporate a human-in-the-loop review process to further refine and validate the dataset. We then fine-tune an open-source MLLM (Qwen2.5VL [wang2024qwen2]) using Struct2D-Set. Although the fine-tuned model is evaluated under noisy 3D perception conditions, it achieves strong performance across multiple spatial reasoning benchmarks, including 3D question answering [yang2024thinking, azuma2022scanqa, ma2022sqa3d], spatial captioning [chen2021scan2cap], and object grounding [chen2020scanrefer, Multi3DRefer], demonstrating the practicality and robustness of our approach.
@@
  • We propose a perception-guided 2D prompting strategy, Struct2D Prompting, and conduct a detailed zero-shot analysis that reveals MLLMs’ ability to perform 3D spatial reasoning from structured 2D inputs alone.
@@
  • We fine-tune an open-source MLLM to achieve competitive performance across several spatial reasoning benchmarks, validating the real-world applicability of our framework.
@@ 2 Related Work
3D Spatial Reasoning with MLLMs. Developing real-world embodied agents requires equipping Multimodal Large Language Models (MLLMs) with robust 3D spatial reasoning abilities [cai2024spatialbot, chen2024spatialvlm, cheng2024spatialrgpt, li2024topviewrs, liu2024coarse, yang2024v, zhu2024llava]. Recent efforts have explored spatial understanding through language [momennejad2023evaluating, rozanova2021grounding, wu2024visualization], static 2D images [ramakrishnan2024does, tang2024sparkle, yang2023set, marsili2025visual, ma2025spatialreasoner], or videos [qi2025gpt4scene, yang2024thinking, liao2025improved, gu2025phyworldbench]. Our work builds upon the video-input setting, but diverges by enabling spatial reasoning in MLLMs using only structured 2D inputs—BEV images, object marks, and metadata—without relying on explicit 3D encoder / representations at inference.
@@
Instruction Tuning for 3D Spatial Reasoning. Recent work [Robin3D, li2023m3dbench, li20243dmit, chen2024ll3da, liao2025improved] has explored instruction tuning to enhance MLLMs’ capabilities for 3D spatial reasoning, targeting tasks such as 3D visual grounding [achlioptas2020referit3d, chen2020scanrefer, Multi3DRefer], 3D dense captioning [chen2021scan2cap], and 3D question answering [azuma2022scanqa, ma2022sqa3d]. M3DBench [li2023m3dbench] provides region- and scene-level instruction-response pairs for general 3D understanding, while 3DMIT [li20243dmit] focuses on scene-centric instructions. LL3DA [chen2024ll3da] supports interactive planning and reasoning across omni-3D inputs. Robin3D [Robin3D] introduces a 3D LLM trained on diverse instruction-following examples. R1-Zero-VSI [liao2025improved] proposes a video-based instruction tuning dataset and a GRPO-based training method to enhance spatial reasoning in Qwen2-VL, yet its QA pairs involve limited reasoning complexity, cover fewer task types, and yield marginal performance gains. In contrast, we propose Struct2D-Set, a large-scale dataset that enables open-source MLLMs to acquire rich 3D spatial reasoning skills through instruction tuning—using only structured 2D representations, without requiring direct access to 3D point clouds.
@@
3D Point Cloud LLMs. Recent advances in 3D point cloud LLMs enable natural language generation and interaction grounded in 3D geometry by directly processing point clouds as input. These models benefit from the geometric precision and texture richness of point clouds, offering stronger spatial understanding than raw image or video inputs. Prior work has focused on object-level [guo2023point, qi2024shapellm, qi2024gpt4point, xu2024pointllm] and scene-level [chen2024ll3da, fu2024scene, hong20233d, man2024lexicon3d, wang2023chat, huang2023chat] spatial reasoning. However, directly using point cloud features requires additional training and infrastructure, limiting flexibility and scalability in real-world applications.
@@
Prompting LLMs. Despite the rapid scaling of large language models (LLMs)[brown2020language, chowdhery2023palm, touvron2023llama, zhang2022opt, achiam2023gpt, hurst2024gpt, dubey2024llama], their reasoning capabilities remain heavily dependent on effective prompt design. In-context learning[brown2020language, dong2022survey], which conditions models on a few representative examples, has become a widely adopted technique for improving instruction-following behavior. To further enhance reasoning, strategies such as chain-of-thought [wei2022chain] and tree-of-thought [yao2023tree] prompting have been proposed. More recently, Multimodal Large Language Models (MLLMs)[o3, gemini, liu2023visual, zhu2023minigpt, zhang2023llama, dai2023instructblip, li2024multimodal, chen2024internvl, xue2024longvila, lin2023vila, zhang2024longva, zhang2024llavanext-video, li2024llava] have gained prominence for their ability to reason over multiple input modalities. This has led to a surge of research into prompting techniques tailored for MLLMs[yang2023dawn, yang2023set, cheng2024spatialrgpt, cai2024vip, hu2025visual, nasiriany2024pivot, wu2025controlmllm, wu2024dettoolchain, yu2024attention, hu2025leveraging, lin2024draw, lee2024affordance, zhang2024earthmarker, zheng2024tracevla, liu20233daxiesprompts, sun2024layoutvlm, wu2024visual, yang2025magma]. Building on this direction, we propose Struct2D, a structured 2D prompting strategy that enables MLLMs to perform 3D spatial reasoning effectively—without requiring explicit 3D input representations.
@@ 3.1 Struct2D Prompting
Given a video $\mathbf{V}$ as input, an MLLM $\mathcal{F}$ processes a set of $N$ sampled video frames, denoted as $\mathbf{I}=\{I_{1},I_{2},\dots,I_{N}\}$, where each frame $I_{n}$ has dimensions $\mathbb{R}^{H\times W\times 3}$ for $n\in\{1,...,N\}$. Alongside visual input, the MLLM receives a text query of length $l_{i}$, represented as $\mathbf{T}^{\text{in}}=[t^{i}_{1},\dots,t^{i}_{l_{i}}]$. The model then generates a textual response of length $l_{o}$, denoted as $\mathbf{T}^{\text{out}}=[t^{o}_{1},\dots,t^{o}_{l_{o}}]$, formulated as:
@@
However, directly using video frames for spatial reasoning introduces two major limitations: (1) Incomplete perception — Video frames are typically sampled sparsely and from limited viewpoints, which can result in missing critical visual evidence required for spatial understanding. For instance, consider a scene where a chair is tucked partially under a table. If most sampled frames are taken from frontal views or from a standing height, the chair’s presence might be obscured or entirely invisible, leading the model to incorrectly assume there is empty space beneath the table. This limitation becomes more severe in cluttered or occluded environments, where small objects or those blocked by other furniture. (2) Lack of global context — Video frames offer fragmented, egocentric views that often fail to capture the overall structure of the scene. For example, determining whether a lamp is closer to the couch or the bookshelf may be impossible if the two objects never co-occur in the same frame. Without a consistent top-down or holistic representation, the model must rely on spatial memory or reasoning across disjoint perspectives—an ability that remains weak in most MLLMs. This fragmentation also impedes the understanding of traversability (e.g., identifying a clear path from the door to the kitchen) or relational queries (e.g., which chair is directly behind the dining table).
@@
This formulation illustrates how Struct2D leverages 3D perception as an intermediate step to generate informative 2D inputs that preserve spatial structure. Although 3D point clouds are used during preprocessing, they are not directly provided to the MLLM. Instead, they are transformed into BEV images and metadata used for prompting. As a result, the model performs spatial reasoning effectively without requiring explicit 3D representations as input.
@@
Evaluation Setup. We sample questions from VSI-Bench [yang2024thinking], which is designed to evaluate complex spatial reasoning skills. Compared to traditional 3D QA datasets [azuma2022scanqa, ma2022sqa3d], VSI-Bench covers more fine-grained object perception requirements, intricate global spatial relationships, and egocentric-to-allocentric transformations. It also features diverse indoor scene sources and robust evaluation metrics that go beyond rule-based NLP scoring.
Comparison to GPT4Scene Prompting [qi2025gpt4scene]. While GPT4Scene pioneered 2D spatial prompting using BEV images, our Struct2D strategy introduces several key improvements: (1) Filtered object marks tailored to the query improve visual relevance and reduce distraction; (2) Guided metadata prompts provide additional spatial priors; (3) Keyframe selection is optimized using depth-aware 3D projection instead of uniform sampling, making them both fewer and more informative (training drops from 6 to 4 hours).